# SN2023syz and SN2025cbj: Two Type IIn Supernovae Associated with IceCube High-energy Neutrinos

### v1

@@ 3 Model-predicted neutrino numbers from SN 2025cbj and SN 2023syz
where the $R_{\rm{ph}}$ is the photosphere radius at the optical depth $\tau_{\rm{T}}(R_{\rm ph})$ = 1. Then we require the parameter space to satisfy $t_{\rm{rise}}\leq t_{\rm{rise,obs}}\leq 1.5\times t_{\rm{rise}}$ (T. Pitik et al., 2023), namely we allow an error up to 50% on the $t_{\rm{rise}}$ estimated from the model. Furthermore, the $t_{\rm CSM}-t_{\rm bo}$ is required to be larger than the time delay between the arrival dates of alert events and SNe IIn (i.e. $t_{\rm CSM}-t_{\rm bo}\geq$ 38 days for SN2023syz and 61 days for SN2025cbj, respectively). This criterion ensures that the forward shock is still interacting with the CSM at the moment the alert event occurred.
Figure 3: Expected total number of neutrinos from SN 2023syz (left panel) and SN 2025cbj (right panel). We scan the parameters of ($M_{\rm ej}$, $M_{\rm CSM}$) = ($_{\odot}, [120]M[1-20]\,M_{\odot}$; ($M_{\rm ej}$, $M_{\rm CSM}$) = ($_{\odot}, [170]M[1-70]\,M_{\odot}$ ) to show how the expected number relies on these parameters for SN2023syz and SN2025cbj, respectively. The gray shaded region in the figure represents the excluded parameters for which the expected $t_{\rm rise}$ and $t_{\rm{CSM}}-t_{\rm{bo}}$ do not match the observations.
In Figure 3, we can see that the expected number of detected neutrinos depends on the values of $M_{\mathrm{ej}}$ and $M_{\mathrm{CSM}}$. For SN 2023syz, its narrow parameter range is mainly due to the short rise time $t_{\mathrm{rise}}$ of this SN. Within the considered parameter space, using the most optimistic choice of the parameters (i.e., the red part in the parameter space; note that the value of $E_{k}\sim 10^{52}\,\rm erg$ we have adopted for these two sources, especially SN 2023syz, is probably also optimistic), the expected numbers of detected neutrinos can reach $N_{\nu_{\mu}+\bar{\nu}_{\mu}}\sim$ 0.0016 for SN2023syz and $N_{\nu_{\mu}+\bar{\nu}_{\mu}}\sim$0.0014 for SN2025cbj. If we simply assume that other SNe IIn also have a comparable neutrino yield (about 0.0015 per source), then for 163 SNe IIn in our sample the total expected number of detected neutrinos is about $\sim 0.24$. This implies that in the optimistic scenario detecting 1–2 neutrino alerts is possible and our model calculation does not rule out the possibility that the two detected neutrino events originated from the SNe.
@@ 4 Summary
Our analysis identify two SNe-neutrino coincidences: SN2023syz ($z=0.037$) with neutrino event IC231027A (38-day delay after supernova onset) and SN2025cbj ($z=0.0675$) with IC250421A (61-day delay). Both neutrinos arrived within the localization uncertainty of their respective supernovae. Monte Carlo simulations estimate that the chance probability of observing $\geq 2$ such coincidences in our sample is only $\sim$0.67%, suggesting these associations may be physical real rather than by chance, though not yet statistically conclusive. To assess the physical plausibility of these associations, we perform model calculations of neutrino emission from these two sources. Using a neutrino production model involving ejecta evolution, shock acceleration, and $pp$ interactions, we compute the expected muon neutrino event rates at IceCube, accounting for CSM properties (mass $M_{\rm CSM}$, radius $R_{\rm CSM}$), ejecta kinetic energy ($E_{k}$) and mass ($M_{\rm ej}$), and observational constraints from the rise time $t_{\rm rise}$ of the optical light curves. We find that, for the most optimistic parameters investigated in this work (the red regions in Figure 3), the predicted numbers of muon neutrino events detectable by IceCube reach $N_{\nu_{\mu}+\bar{\nu}_{\mu}}\sim 0.0016$ events for SN 2023syz and $N_{\nu_{\mu}+\bar{\nu}_{\mu}}\sim 0.0014$ events for SN 2025cbj, respectively. This implies that totally detecting 1-2 neutrino alerts from the 163 SNe IIn is possible. The model-expected number of neutrino events from these SNe IIn can accommodate the actual detections.
This work provides supportive (though not conclusive) evidence that interacting supernovae, particularly SNe IIn, could be sources of high-energy neutrinos. Though the limited sample size precludes definitive claims, the relatively low chance probability and model compatibility suggest that SNe IIn may contribute to the high-energy neutrino sky. Future observations with enhanced neutrino sensitivity and more powerful transient surveys (e.g., Vera Rubin Observatory (Ž. Ivezić et al., 2019; K. Bricman & A. Gomboc, 2020)) will help confirm these associations and quantify the role of SNe IIn in cosmic neutrino production.

### v3

@@ 3 Model-predicted neutrino counts from SN 2025cbj and SN 2023syz
where the $R_{\rm{ph}}$ is the photosphere radius at the optical depth $\tau_{\rm{T}}(R_{\rm ph})$ = 1. Then we require the parameter space to satisfy $t_{\rm{rise}}\leq t_{\rm{rise,obs}}\leq 1.5\,t_{\rm{rise}}$ (T. Pitik et al., 2023), namely we allow an error up to 50% on the $t_{\rm{rise}}$ estimated from the model. Furthermore, the $t_{\rm CSM}-t_{\rm bo}$ is required to be larger than the time delay between the arrival dates of alert events and SNe IIn (i.e. $t_{\rm CSM}-t_{\rm bo}\geq$ 38 days for SN2023syz and 61 days for SN2025cbj, respectively). This criterion ensures that the forward shock is still interacting with the CSM at the moment the alert event occurred.
(a) 1-100 TeV for SN2023syz
(b) 0.1-10 PeV for SN2023syz
(c) 1-100 TeV for SN2025cbj
(d) 0.1-10 PeV for SN2025cbj
Figure 3: Expected total number of neutrinos, integrated over different energy ranges, from SN 2023syz (upper panels) and SN 2025cbj (lower panels). We scan the parameters of ($M_{\rm ej}$, $M_{\rm CSM}$) = ($_{\odot}, [120]M[1-20]\,M_{\odot}$); ($M_{\rm ej}$, $M_{\rm CSM}$) = ($_{\odot}, [170]M[1-70]\,M_{\odot}$) to show how the expected number relies on these parameters for SN2023syz and SN2025cbj, respectively. The gray shaded region in the figure represents the excluded parameters for which the expected $t_{\rm rise}$ and $t_{\rm{CSM}}-t_{\rm{bo}}$ do not match the observations.
[…]
Next, we estimate the expected detection probability of neutrinos from the entire SN IIn population. We assume that the expected neutrino number for each source is related to the IceCube effective area $A_{\rm eff}$ (which is related to the DEC and energy of alerts) and the supernova’s peak flux ($F_{\rm peak}$) or luminosity distance ($1/d_{L}^{2}$):
$N_{\rm expect}\propto\int_{E_{\nu,1}}^{E_{\nu,2}}{A_{\rm eff}(E_{\nu},{\rm DEC})}\times X\,dE_{\nu}.$ (10)
with $X$ being $F_{\rm peak}$ or $1/d_{L}^{2}$. The reason for us to adopt peak flux and distance as the weighting factors in the calculation is that the former can serve as a proxy for the SN’s kinetic energy, while the latter assumes that the distribution of kinetic energies is the same for SNe at different distances. Using the mean value of $N_{\nu_{\mu}+\bar{\nu}_{\mu}}$ obtained from SN2023syz and SN2025cbj (approximately $2.5\times 10^{-4}$ per source, since the expected neutrino counts from SN2023syz is negligible), we estimate that in the optimistic scenario, the total expected number of neutrinos from 163 SNe IIn is $\sim$0.06 (0.11) when $X$ is $F_{\rm peak}$ ($1/d_{L}^{2}$), with the contributions from the energy ranges 1–100 TeV and 0.1–10 PeV approximately 0.03 (0.04) and 0.03 (0.07), respectively. Our model calculations show that, under optimistic parameter choices, there is a certain probability ($\sim 5\%$) to detect one neutrino alert event. Therefore, the observed neutrino event in the direction of SN2025cbj originating from this supernova is still possible.
@@ 4 Summary
Our analysis identify two SNe-neutrino coincidences: SN2023syz ($z=0.037$) with neutrino event IC231027A (38-day delay after supernova onset) and SN2025cbj ($z=0.0675$) with IC250421A (61-day delay). The later is also reported in S. Garrappa et al. (2025). Both neutrinos arrived within the localization uncertainty of their respective supernovae. Monte Carlo simulations estimate that the chance probability of observing $\geq 2$ such coincidences in our sample is only $\sim$0.67%, suggesting these associations may be physical real rather than by chance, though not yet statistically conclusive. To assess the physical plausibility of these associations, we perform model calculations of neutrino emission from these two sources. Using a neutrino production model involving ejecta evolution, shock acceleration, and $pp$ interactions, we compute the expected muon neutrino event rates at IceCube, accounting for CSM properties (mass $M_{\rm CSM}$, radius $R_{\rm CSM}$), ejecta kinetic energy ($E_{k}$) and mass ($M_{\rm ej}$), and observational constraints from the rise time $t_{\rm rise}$ of the optical light curves. For SN2025cbj, the predictions based on the most optimistic parameter sets explored in this work (the yellow regions in Figure 3) indicate that the number of detectable muon neutrino events, $N_{\nu_{\mu}+\bar{\nu}_{\mu}}$, is $\sim 1.8\times 10^{-4}$ in the 1–100 TeV band and $\sim 3.1\times 10^{-4}$ in the 0.1–10 PeV band. The expected number of neutrino events from SN2023syz is estimated to be at the $10^{-7}$ level and is negligible, making its association with IC231027A likely a chance coincidence rather than physically real. When considering the most optimistic parameter set, our calculation shows that the total expected number of neutrinos detectable from the sample of 163 Type IIn supernovae is 0.06 $\sim$ 0.11 (depending on how different SNe are weighted in the calculation), and the detection of one neutrino coincidence is still possible. It should also be noted that when more common or typical (i.e., not that optimistic) SN parameters are adopted, the expected number of neutrinos may significantly decrease, making the detection of neutrinos from SNe much less likely.
This work provides important insights into assessing whether interacting supernovae, particularly SNe IIn, could be potential sources of high-energy neutrinos. Our analysis shows a relatively low probability of chance coincidence. Although model calculations indicate a low expected number of neutrino events, detection remains possible under suitable parameters, and the physical association between SN2025cbj and IC250421A cannot yet be ruled out. Type II superluminous supernovae (SLSNe-II) exhibit higher emission power and ejecta kinetic energy. If the detected association were physically real, one would expect more neutrino events to be observed from SLSNe. However, the current absence of any observed associations between SLSNe and neutrino alerts poses a challenge to such a model scenario. We leave the investigation of SLSNe to future work. Future observations with enhanced neutrino sensitivity and more powerful transient surveys (e.g., Vera Rubin Observatory (Ž. Ivezić et al., 2019; K. Bricman & A. Gomboc, 2020)) will help confirm or rule out the associations between SNe IIn and neutrinos definitively and quantify the role of SNe IIn in cosmic neutrino production.
@@
Appendix A KS-test in the Monte Carlo simulation
In the Monte Carlo simulation, we apply a KS test to the distributions of $\rm RA$ and $\rm DEC$ for the simulated sources to ensure that the mock samples preserve the distribution characteristics of the real sources. The comparison of spatial distribution between mock sources and real ones after applying the KS test is presented in Figure 4 for a demonstration. The mock source catalog generated in this way has almost identical spatial distribution as the real catalog.
Figure 4: A comparison of the RA and DEC distributions between the real SN IIn catalog and one realization of the mock catalogs.
Appendix B Fermi-LAT data analysis around IC231027A
The gamma-ray sources of the fourth Fermi catalog (4FGL-DR4, for Data Release 4) [^9] https://fermi.gsfc.nasa.gov/ssc/data/access/lat/14yr_catalog/(S. Abdollahi et al., 2020, 2022) has two sources located nearby the IC231027A ($0.3^{\circ}$), as reported by (S. Buson et al., 2023). We take the Fermi-LAT data to analyze these two sources. Centered on the best fit position of IC231027A, we use Pass 8 SOURCE events (evclass=128, evtype=3) with energies ranging from 100 MeV to 1 TeV, and select a $15^{\circ}\times 15^{\circ}$ Region of Interest (ROI). The P8R3_SOURCE_V3 of LAT instrument response functions are adopted. The Galactic diffuse and isotropic component use the file of gll_iem_v07.fits and iso_P8R3_SOURCE_V3_v1.txt, respectively. We exclude events with zenith angle $z_{\rm max}<90^{\circ}$ to reduce the contamination produced by the Earth’s limb. We first make a TS map integrating $\pm$ 120 days of arrival date of IC231027A for which find gamma-ray counts nearby the IC231027A. The significant gamma-ray counts are not found during the time span, as shown in figure 5. Then, for these two 4FGL-DR4 sources as reported in S. Buson et al. (2023), we perform a likelihood analysis, with spatial bins of 0.1 and 20 logarithmically-space bins for energy. We adopt the spatial model and spectrum as recorded in the model file of gll_psc_v32.xml.
We present the light curves for these two 4FGL-DR4 sources in figure 6, where upper limits are shwon for time bins with TS<9. The light curves span a period from 90 days before to 120 days after the arrival date of IC231027A, with a bin size of 30 days. The obvious correlation is not found between Fermi-LAT light curves and arrival time of IC231027A. Therefore, we conclude that these two Fermi-LAT sources are unlikely to be associated with IC231027A.
Figure 5: This figure is the Fermi-LAT counts map in the 100 MeV - 1 TeV energy range, showing the integrated search in the period of $\pm$ 120 days of the arrival date of IC231027A. The position of IC231027A is marked by the red ’X’ with the 90% positional uncertainty denoting by the red circle (only statistical error). The orange star is the position of SN2023syz. The white circle and plus sign are the positions of two 4FGL sources within the error circle of IC231027A (i.e. 4FGL J1749.6+4716 and 4FGL J1747.9+4704), away from the best fit position of IC231027A 0.16 deg and 0.29 deg, respectively.
Figure 6: Fermi-LAT light curves of the two 4FGL-DR4 sources (i.e. 4FGL J1749.6+4716 and 4FGL J1747.9+4704) near the best position of IC231027A ($0.3^{\circ}$). We integrate the energies from 100 MeV to 1 TeV. The time spans 90 days before the arrival date of IC231027A and 120 days after it with a 30-day bin. We plot upper limits for points with ${\rm TS}<9$. We don’t find any significant gamma-ray signal around the arrival date of IC231027A.
Appendix C The choice of kinetic energy
Given the range of radiative efficiency, the $E_{k}$ could be evaluated an approximate range from the total radiated energy ($E_{\rm rad}$), following the approach of T. Pitik et al. (2022, 2023). However, the lack of multi-wavelength observations for these two sources restricts the direct evaluation on $E_{\rm rad}$. Therefore, we utilize a blackbody model applied to the ZTF-BTS data to roughly estimate their $E_{\rm rad}$. The blackbody model [^10] The blackbody model is described as https://heasarc.gsfc.nasa.gov/docs/software/xspec/manual/node136.html is characterized by the temperature ($kT_{\rm bb}$, where $k$ is the boltzmann constant) and the blackbody luminosity ($L_{\rm bb}$). We constrain the values of these two parameters by fitting the data points obtained from ZTF-BTS observations at six epochs. The continuous light curves of SN2023syz and SN2025cbj are generated using linear interpolation of the data in the $r$ and $g$ band, as show in figure 7. For instance, the peak luminosity of SN2025cbj could be described by the blackbody model with parameters ($kT_{\rm bb}$, $L_{\rm bb}$) = (3 eV , 2.3 $\times$ 1044 erg/s), as shown in figure 8.
We fit these light curves at six epochs as follows. The first epoch is the initial observation. The second epoch is the peak luminosity. Then the third epoch is where the the $r$-band luminosity has the same as the $g$-band one. The fourth and the fifth epochs are selected from periods where the $r$-band luminosity exceeds the $g$-band luminosity. The final epoch corresponds to the last data point. We present these epochs by the dashed black lines in figure 7. However, note that the parameters constraints are limited by the availability of only two bands from ZTF-BTS.
Figure 7: Left panel: the ZTF-BTS light curve of SN2023syz. The red and green points are the $r$ band and $g$ band luminosity, respectively. The dashed black lines are the six epochs fitted by the blackbody model (see details in Appendix C). Right panel: the same as the left panel, but for SN2025cbj.
The estimated parameter ranges for the blackbody components are as follows. For SN2023syz, the $kT$ and $L_{\rm bb}$ are the 0.5-3.1 eV and (0.41-5.27) $\times$ 1043 erg/s, respectively. For SN2025cbj, the $kT$ and $L_{\rm bb}$ are the 0.5-4.3 eV and (0.06-2.52) $\times$ 1044 erg/s, respectively. We can estimate $E_{k}$ for SN2023syz and SN2025cbj based on assumed mean blackbody temperature of $kT\sim 2$ eV and $kT\sim 3$ eV, respectively. Incorporating the uncertainty in radiative efficiency ($E_{\rm rad}$ = 0.1 $\sim$ 0.8 $E_{k}$(T. Pitik et al., 2023), the derived $E_{k}$ ranges are (1-8) $\times$ 1050 erg for SN2023syz and (0.3-3) $\times$ 1052 erg for SN2025cbj.
Furthermore, we also search for X-ray counterparts to SN2023syz and SN2025cbj based on the Living Swift XRT Point Source Catalog (LSXPS) [^11] https://www.swift.ac.uk/LSXPS/ covering the energy range of 0.3-10 keV (P. A. Evans et al., 2022). This search yields no detections within $0.05^{\circ}$ of their positions. This lack of detected X-ray emission is consistent with the negligible blackbody flux estimated from their $E_{k}$ range and $kT$ adopted in this work.
Figure 8: Fit of the observational data of SN2025cbj with a blackbody spectral model. The black line corresponds to the model parameters of $(kT,L_{\rm bb})\sim(3\,{\rm eV},2.3\times 10^{44}\,{\rm erg/s})$. The red and green points represent the $r$-band and $g$-band data, respectively, from the second epoch (i.e., the peak epoch) of SN2025cbj.